Sullivak24:

The lemonade comes in 100% juice concentrate, but you only serve it as 70% solution. Unfortunately, one batch got overwatered, so you have 4 quarts of 50% solution. How much 100% concentrate do you need to add in order to get the 70% solution? How much of the 70% solution will you have?

Vocaloid:

let x be the number of quarts of the 100% solution added right now we have 4 * (0.50) = 2 quarts of juice inside 4 quarts of solution (juice)/(total solution) = (0.7) as desired when you add x quarts of 100% concentrate, you are adding x quart of pure juice and adding x qrt of volume to the total solution so (2 + x)/(4+x) = 0.7, solve for x

Vocaloid:

hope I did this right, if someone could double check me that would be really appreciated

Gdeinward:

Seems logical and accurate to me. I am currently attempting to solve said equation.

Gdeinward:

I got x = 8/3